A particle is falling from height h calculate velocity when it strike to ground.
1
Expert's answer
2020-11-10T06:59:47-0500

Solution

When a particle is falling free from height h then it's initial velocity is u=0

Here acceleration is gravitation acceleration g, let velocity is v at ground


Then using third equation of motion


"v^2=u^2+2as"

"v^2=0+2gh"

"v=\sqrt{2gh}"